OpenStudy (linn99123):

The area of the base of a prism is 55 mm2. The perimeter of the base is 30 mm. The height of the prism is 7 mm. What is the surface area of the prism? A. 445 mm2 B. 320 mm2 C. 265 mm2 D. 92 mm2 I picked C

OpenStudy (johnweldon1993):

Well lets see \[\large SA = \text{Perimeter of base} \times H + 2(\text{Area of base})\] \[\large SA = 30 \times 7 + 2(55)\] \[\large SA = 210 + 110\] \[\large SA = 320\]